Today you purchase a coupon bond that pays an annual interest, has a par value of $1,000, matures in six years, has a coupon rate of 10%, and has a yield to maturity of 8%. One year later, you sell the bond after receiving the first interest payment and the bond's yield to maturity had changed to 7%. Your annual total rate of return on holding the bond for that year is ?
Answer #1

Price=Par value*Coupon rate/yield*(1-1/(1+yield)^t)+Par value/(1+yield)^t

Purchase Price=1000*10%/8%*(1-1/1.08^6)+1000/1.08^6=1092.457593

Selling Price=1000*10%/7%*(1-1/1.07^5)+1000/1.07^5=1123.005923

Total rate of return=(1123.005923+1000*10%)/1092.457593-1=11.95%